A 4.0 kg ball is swung at a constant speed in a vertical circle of radius 2.5 m. If the ball completes one revolution in 3.0 s, what is the tension in the rope at the top of the circle?


1
Expert's answer
2021-07-25T09:16:02-0400
"T=m\frac{v^2}{r}-mg\\T=m\frac{(2\pi r)^2}{rT^2}-mg
\\T=m\frac{(2\pi )^2r}{T^2}-mg\\T=4\frac{(2\pi )^2(2.5)}{3^2}-4(9.8)=4.7\ N"